Show that

f(x)=(1/2)^(x+1) , x = 0, 1, 2, 3, 4, 5,...

is a valid PMF for a discrete random variable. Also find

out its CDF.


1
Expert's answer
2022-06-06T16:51:44-0400

a.


"\displaystyle\sum_{x=0}^{\infin}(\dfrac{1}{2})^{x+1}=\dfrac{1/2}{1-1/2}=1, True"

Therefore "f(x)(\dfrac{1}{2})^{x+1}, x=0,1,2,3,4,5,..." is a valid PMF for a discrete random variable.


b.


"F(x)=\displaystyle\sum_{y\le x}^{x}f(y)=\displaystyle\sum_{y=0}^{x}(\dfrac{1}{2})^{y+1}=\dfrac{1}{2}\displaystyle\sum_{y=0}^{x}(\dfrac{1}{2})^{y}"

"=\dfrac{1}{2}(\dfrac{1-(1/2)^{x+1}}{1-1/2})=1-(\dfrac{1}{2})^{x+1}"

"F(x)= \begin{cases}
 0&x<0 \\
 1-(\dfrac{1}{2})^{x+1}&x=0,1,2,3,4,5,...
\end{cases}"